If two opposite sides of a square are increased by 6 meters and the other sides are decreases by 4 meters, the area of the rectangle that is formed is 96 square meters. Find the area of the original square.
I've gotten as far as finding x at least I think 10, -12...but I don't know where to go from there. :(

Answer by checkley71(8403)

(x+6)(x-4)=96
x^2+6x-4x-24=96
x^2+2x-24-96=0
x^2+2x-120=0
(x+12)(x-10)=0
x-10=0
x=10 meters is the length of the square sides.
PROOF:
(10+6)(10-4)=96
16*6=96
96=96